# Shot-by-Shot: Film-Grammar-Aware Training-Free Audio Description Generation

### v1

@@ 1 Introduction
In this paper, we develop a training-free AD generation approach. Inspired by AutoAD-Zero [90], as illustrated in Fig. 1, we adopt a two-stage pipeline in which a VideoLLM generates dense text descriptions in the first stage, followed by an LLM that produces the final AD outputs from this text. We improve the effectiveness of training-free approaches by incorporating two key improvements based on: (i) shot-based temporal context; and (ii) the shot scale and thread structure, that are both crucial to cinematic composition and understanding.
@@ 3 Training-free AD generation framework
In Stage I, we employ a VideoLLM to generate a dense description for the clip based on instructions :
@@ 3.2 Leveraging thread structure
Thread structure prediction. We develop a training-free method to predict thread clustering. The problem here is to determine if two shots correspond to the same viewpoint or not. Specifically, given two shots (Shot and a later Shot ), we compare the last frame of Shot () with the first frame of Shot (), extracting their DINO features as and .
@@ 3.3 Leveraging shot scale information
Shot scale classification. We first build a shot scale classification network by fine-tuning a pre-trained DINO model. For all current shots, we classify their shot scales into one of five classes, represented by values –, as shown in Fig. 4 (top). We then compute their average to obtain the effective shot scale .
@@ 4 Action score
Action score computation. The action score features a combination of semantic-based and verb-matching-based components, as illustrated in the middle and right column of Fig. 5, respectively.
@@ 5.3 Thread structure prediction
Tab. 1 verifies the choice of DINOv2-g14 as the feature extractor for frame pair comparison. Compared to abstract CLS tokens, dense spatial matching achieves higher AP in frame pair relationship prediction and higher WCP in thread clustering. Additionally, we observe that using the temporally closest frames from two shots (“Side”) leads to a noticeable performance improvement. This can be attributed to the continuous story flow across shots within the same thread.
@@ 5.4 Shot scale classification
Since previous state-of-the-art methods on shot scale classification are not open-sourced, we develop a new network by fine-tuning DINOv2, achieving superior performance over prior approaches that rely on additional optical flow or SAM-based mask inputs, as demonstrated in Tab. 2.
@@ Acknowledgements.
This research is supported by the UK EPSRC Programme Grant Visual AI (EP/T028572/1), a Royal Society Research Professorship RPR1191132, a Clarendon Scholarship, and ANR-21-CE23-0003-01 CorVis.
@@ Appendix A Implementation details
[…]
@@ Appendix B Evaluation metrics for AD generation
LLM-AD-Eval [29] employs LLM agents (LLaMA3-8B [53] LLaMA2-7B [19]) as evaluators to compare ground truth ADs with predictions, generating a matching score ranging from 1 (lowest) to 5 (highest).
@@
| Thread structure | Stage I VideoLLM | TV-AD subset | ||
| CIDEr | CRITIC | Action | ||
| ✗ | Qwen2.5-VL-7B [8] | |||
| Qwen2.5-VL-7B [8] |
1.2 |
0.6 |
0.0 |
[…]
| VideoLLaMA3-7B [97] |
0.6 |
1.0 |
0.8 |
[…]
| InternVL2.5-8B [18] |
1.2 |
1.6 |
1.3 |
[…]
| Qwen2-VL-7B [83] |
1.9 |
0.7 |
0.3 |
@@ Appendix C Additional experimental results
Different VideoLLMs for Stage I. Tab. A1 compares our AD generation performance using different open-source VideoLLMs in Stage I, validating our choice of Qwen2-VL-7B as the default model.
@@
Thread structure injection for different VideoLLMs. We investigate how injecting thread structure into different open-source VideoLLMs affects AD generation. Specifically, we evaluate performance on TV-AD, which contains a large proportion of thread-structured video clips. As shown in Tab. A2, incorporating thread information leads to general performance boosts across various VideoLLMs.
Repeated AD generation. Tab. A3 presents the results of repeated AD generation experiments, demonstrating consistency across different runs. In particular, the CRITIC results exhibit the highest variance, followed by CIDEr, while the action score remains relatively stable across repeated experiments.

### v2

@@ 1 Introduction
In this paper, we develop a training-free AD generation approach. Inspired by AutoAD-Zero [92], as illustrated in Fig. 1, we adopt a two-stage pipeline in which a VideoLLM generates dense text descriptions in the first stage, followed by an LLM that produces the final AD outputs from this text. We improve the effectiveness of training-free approaches by incorporating two key improvements based on: (i) shot-based temporal context; and (ii) the shot scale and thread structure, that are both crucial to cinematic composition and understanding. Previous methods could describe human-object interactions or human-human interactions, such as looks [52], within a frame or shot. With the improvements we introduce, the generated AD also includes such interactions when they are implied by the shot structure.
@@ 3 Training-free AD generation framework
In Stage I, we employ a VideoLLM that takes a sequence of frames (from multiple shots) as input and generates a dense description , guided by instructions :
@@ 3.2 Leveraging thread structure
Thread structure prediction. We develop a training-free method to predict thread clustering. The problem here is to determine if two shots correspond to the same viewpoint or not. Specifically, given two shots (Shot and a later Shot ), we compare the last frame of Shot () with the first frame of Shot (), extracting their DINOv2 [61] features as and .
@@ 3.3 Leveraging shot scale information
Shot scale classification. We first build a shot scale classification network by fine-tuning a pre-trained DINOv2 [61] model. For all current shots, we classify their shot scales into one of five classes, represented by values –, as shown in Fig. 4 (top). We then compute their average to obtain the effective shot scale .
@@ 4 Action score
Action score computation. The action score features a combination of semantic-based and verb-matching-based components, as illustrated in the middle and right columns of Fig. 5, respectively.
@@ 5.3 Thread structure prediction
Tab. 1 verifies the choice of DINOv2-g14 [61] as the feature extractor for frame pair comparison. Compared to abstract CLS tokens, dense spatial matching achieves higher AP in frame pair relationship prediction and higher WCP in thread clustering. Additionally, we observe that using the temporally closest frames from two shots (“Side”) leads to a noticeable performance improvement. This can be attributed to the continuous story flow across shots within the same thread.
@@ 5.4 Shot scale classification
Since previous state-of-the-art methods on shot scale classification are not open-sourced, we develop a new network by fine-tuning DINOv2 [61], achieving superior performance over prior approaches that rely on additional optical flow or SAM-based mask inputs, as demonstrated in Tab. 2.
@@ Acknowledgements.
This research is supported by the UK EPSRC Programme Grant Visual AI (EP/T028572/1), a Royal Society Research Professorship RSRPR241003, a Clarendon Scholarship, and ANR-21-CE23-0003-01 CorVis.
@@ Appendix A Implementation details
| Qwen2-VL-7B [85] | Gemma3-12B [78] | ||||||||
| Qwen2-VL-7B [85] | Qwen3-8B [96] | ||||||||
| Qwen2-VL-7B [85] | LLaMA3-8B [53] | ||||||||
| gray GPT-4o [60] | GPT-4o [60] | ||||||||
@@ Appendix B Evaluation metrics for AD generation
LLM-AD-Eval [28] employs LLM agents (LLaMA2-7B [79] LLaMA3-8B [53]) as evaluators to compare ground truth ADs with predictions, generating a matching score ranging from 1 (lowest) to 5 (highest).
@@ Appendix C Additional experimental results
Different VideoLLMs for Stage I. Appendix A compares our AD generation performance using different open-source VideoLLMs in Stage I, validating our choice of Qwen2-VL-7B as the default model.
@@
Different LLMs for Stage II. Appendix A also compares different options for the Stage II LLM. Overall, the choice of LLM has a relatively minor impact compared to the Stage I VideoLLM. The default LLaMA3-8B achieves overall strong performance.
Thread structure injection for different VideoLLMs. We investigate how injecting thread structure into different open-source VideoLLMs affects AD generation. Specifically, we evaluate performance on TV-AD, which contains a large proportion of thread-structured video clips. As shown in Appendix B, incorporating thread information leads to general performance boosts across various VideoLLMs.
Repeated AD generation. Tab. A3 reports results from five independent runs of our two-stage AD generation pipeline, each adopting a different random seed. The results are largely consistent across runs, indicating the stability of the AD generation process. In particular, the CRITIC results exhibit the highest variance, followed by CIDEr, while the action score remains relatively stable across repeated experiments.
@@
| Method | Shot Partition | Film Grammar | Stage I VideoLLM | Stage II LLM | Total |
| AutoAD-Zero | s | s | s | ||
| Ours | s | s | s | s | s |
Inference time analysis. We report the inference time per AD in Tab. A4. Our added components (e.g. shot partitioning, film grammar prediction) incur minimal overhead. The main cost increase arises from sampling more contextual frames for Stage I inputs. Overall, our method maintains reasonable efficiency—given averaged input clip duration (including contextual shots) is s—while delivering clear performance gains.
Investigation on factor–scale correlations. To verify our assumptions about the strong correlations between descriptive factors in ADs and shot scales (as mentioned in Sec. 3.3), we used GPT-4o to extract key elements (e.g. environments) mentioned in GT ADs from movies, and analysed their correlation with shot scales. As shown in Fig. A1, facial expressions predominantly occur in (extreme) close-up shots, while environmental cues align with full/long shots. These observations align with the assumption and generalise across real-world movies.